OpenStudy (anonymous):

algebra unit 7 lesson 10 assessment question 2 Which is an equation of the given line in point-slope form? A. B. C. D.

OpenStudy (anonymous):

is there more info?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

a) y-1=6/7 (x-3) B)y+3=7/6(x+1) C)y+1=6/5(x+3) D)y+1=6/7(x+3)

OpenStudy (anonymous):

a

OpenStudy (anonymous):

and the two points on the graph are (3,1) ad (1/2, 2)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

thank you!:)

OpenStudy (anonymous):

medal please :)